# Shift Happens: Mixture of Experts based Continual Adaptation in Federated Learning

### v1

@@ Abstract.
Federated Learning (FL) enables collaborative model training across decentralized clients without sharing raw data, yet faces significant challenges in real-world settings where client data distributions evolve dynamically over time. This paper tackles the critical problem of covariate and label shifts in streaming FL environments, where non-stationary data distributions degrade model performance and require adaptive middleware solutions. We introduce ShiftEx  a shift-aware mixture of experts framework that dynamically creates and trains specialized global models in response to detected distribution shifts using Maximum Mean Discrepancy for covariate shifts. The framework employs a latent memory mechanism for expert reuse and implements facility location-based optimization to jointly minimize covariate mismatch, expert creation costs, and label imbalance. Through theoretical analysis and comprehensive experiments on benchmark datasets, we demonstrate 5.5-12.9 percentage point accuracy improvements and 22-95% faster adaptation compared to state-of-the-art FL baselines across diverse shift scenarios. The proposed approach offers a scalable, privacy-preserving middleware solution for FL systems operating in non-stationary, real-world conditions while minimizing communication and computational overhead.
@@ 1. Introduction
A critical challenge in streaming FL environments, is that the data from the parties will change and evolve over time. Consider the example shown in Figure 1 depicting satellite imagery to study land-use patterns. We observe differences in the images based on the changing weather conditions (e.g.) image clarity deteriorates in the winter when there is snow and frost. Hence, as new data arrives, both the input distribution $P(X)$ and label distribution $P(Y)$ can shift  (Zhao et al., 2018). These distributional shifts, known as covariate shift and label shift respectively, pose significant challenges for traditional FL approaches that maintain a single global model (Yoon et al., 2021).
@@ 3.2. ShiftEx Architecture
Each party maintains a local dataset and runs a stream processing engine (e.g., Apache Kafka (Garg, 2013) or Flink (Carbone et al., 2015)) to collect, ingest, and preprocess incoming data streams, storing them in a local database. This setup resembles standard streaming data platforms. ShiftEx extends this setup by incorporating a shift detection module that computes latent feature representations and label distributions from local data in consecutive windows. Covariate shift is detected using Maximum Mean Discrepancy (MMD) on the latent representations, while label shift is detected via Jensen-Shannon Divergence (JSD) on label histograms. After computing these statistics, parties transmit the aggregated latent representations, label distributions, and the associated MMD and JSD scores to the aggregator.
@@ 4.2. Covariate Shift Detection via Maximum Mean Discrepancy
The key advantage of MMD is that it provides a principled way to detect distributional differences without requiring distributional assumptions or labeled data. It operates directly on feature representations, making it ideal for privacy-preserving covariate shift detection in federated settings. A high MMD value ($>\delta_{\text{cov}}$) indicates a significant change in the feature distribution, signaling covariate shift.
@@ 5.2.2. Expert Assignment via Latent Memory Matching
Rather than training new experts for every shift, the aggregator maintains a latent memory—an exponential moving average of each expert’s embedding signatures—to enable expert reuse. For each covariate cluster $G_{j}$, the aggregator computes the cluster centroid $\bar{P}_{j}(X)$ and checks for matching existing experts:
If $\text{If }\min_{k}\mathrm{MMD}(\bar{P}_{j}(X),\mathcal{M}^{(k)})\leq\epsilon\text{, then assign }G_{j}\text{ to expert }k$
where $\mathcal{M}^{(k)}$ represents the latent memory signature of expert $k$. This matching mechanism enables efficient expert reuse for recurring shift patterns while maintaining privacy (only aggregate embeddings are compared). Importantly, this design also serves as a practical approximation of the MMD term in Equation 5.1. By clustering parties based on latent signatures and matching cluster centroids to expert memory, we approximate the MMD term in 5.1 efficiently, avoiding costly pairwise comparisons across a large number of parties and reducing computation in high-scale federated settings. Additionally, this mechanism helps minimize the $\lambda$-weighted term in Equation 5.1, which penalizes the creation of new experts. By assigning clusters to existing experts when a match is found, the system prevents unnecessary expert creation and training, promoting parameter reuse and system efficiency.
@@ 5.2.3. Expert Update using FLIPS/Small Groups
When a covariate cluster $G_{j}$ matches an existing expert $\theta^{(k)}$ (based on latent memory matching via MMD), the system assigns the $\theta^{(k)}$ expert to the cluster. FL is then performed using only the matched party cohort $G_{j}$, ensuring expert specialization remains aligned with the current distribution. To prevent overfitting to skewed or imbalanced label distributions, we employ FLIPS (Federated Learning using Intelligent Participant Selection) (Bhope et al., 2023), which clusters parties based on their local label distributions and selects participants equitably from each cluster. This ensures that the aggregated data used in training reflects a balanced label distribution. This label-aware sampling strategy also helps approximate the $\mu$ term in Equation 5.1. By ensuring that the selected cohort’s aggregate label distribution closely matches the i.i.d. reference distribution, the Jensen-Shannon Divergence is reduced, driving the term toward zero and promoting more stable and generalizable expert training. For smaller covariate clusters $|G_{j}|\leq\gamma$, no Federated Training is performed and the parties are signaled to perform local training or transfer learning with their local datasets on the selected expert.
@@ 5.3. TEE for Privacy Preservation
FL, in practice, only guarantees that raw data never leaves the party device. However, FL systems frequently exchange intermediate artifacts such as model updates, feature embeddings, and drift statistics, which can inadvertently leak private information through inference or reconstruction attacks (Hu et al., 2022; DLG git repository, 2020; Zhao et al., 2020; Geiping et al., 2020; Yin et al., 2021; Zhu et al., 2019). These vulnerabilities highlight the need for additional mechanisms on top of standard FL pipelines to ensure end-to-end privacy guarantees. To mitigate these risks, we augment ShiftEx with Trusted Execution Environments (TEEs) like Intel SGX (int, [n. d.]) or AMD SEV (amd, [n. d.]) on the aggregator side, which offers hardware-level isolation for secure and verifiable computation. TEEs ensure that sensitive operations such as drift detection, clustering, and expert model updates are executed in a secure enclave, where the contents are protected from both the central aggregator and potential adversaries. In our design, parties compute and locally store $d$-dimensional feature embeddings using pre-trained encoders. These embeddings are transmitted to the aggregator encrypted within a TEE enclave, where distributional shift detection (via MMD) and clustering operations are securely performed. Since no raw data or labels are exposed outside the enclave, confidentiality is guaranteed. Expert model updates can also be securely aggregated within TEEs to prevent inference attacks on local model parameters. While TEEs introduce minor computation and memory overhead (5 % for AMD SEV) (Bhope et al., 2023) due to enclave initialization and secure memory constraints, they offer a practical trade-off between privacy and performance.
@@ 6. Experimental Strategy
Datasets: We evaluate our framework on five datasets designed to capture a range of realistic covariate and label shifts in FL scenarios. FMoW (Functional Map of the World) (Christie et al., 2018) is a satellite imagery dataset containing over a million images labeled with functional use categories (e.g., hospital, airport, school) across diverse geographical regions and time periods. We select a subset of 10 labels (Jothimurugesan et al., 2023) for our experiments. Next, we use Tiny-ImageNet-C (Hendrycks and Dietterich, 2019) that extends Tiny-ImageNet by introducing 15 corruption types (e.g., fog, snow, brightness, frost) at 5 severity levels. CIFAR-10-C(Hendrycks and Dietterich, 2019) applies similar corruptions to CIFAR-10, supporting controlled small-scale evaluations. We then use 2 benchmark datasets widely used in FL literature, FEMNIST(Caldas et al., 2019) and Fashion-MNIST(Xiao et al., 2017). FEMNIST, a federated variant of EMNIST, includes handwritten characters from hundreds of users, each contributing uniquely styled samples. Fashion-MNIST is a dataset of images of clothing items that are commonly used as a benchmark for machine learning models. We simulate 200 parties for all datasets except FMoW, for which we use 50 parties. Using 200 parties allows us to capture fine-grained heterogeneity in party data distributions, making it ideal for evaluating adaptability under diverse, non-IID conditions. In contrast, FMoW contains rich, high-resolution satellite imagery where even a small number of parties can exhibit significant covariate variation. Reducing the party count to 50 ensures that each party has access to a sufficient volume of diverse samples to support stable shift (naturally occurring) detection. Additionally, we pick 20% of the participant pool in each round of FL.
@@
Models: For each dataset, we use a model architecture suitable for its scale and complexity. We apply LeNet-5 (Lecun et al., 1998) for FEMNIST and FashionMNIST, DenseNet-121 (Huang et al., 2018) for FMoW, ResNet-18 (He et al., 2016) for CIFAR-10-C, and ResNet-50 (He et al., 2016) for Tiny-ImageNet-C. These choices reflect standard baselines in the literature. For detecting covariate shifts, we leverage encoder-based representations derived from the internal layers of trained party-side models. Specifically, we extract feature embeddings from the penultimate (pre-logit) layer of each model architecture—ResNet, DenseNet, and LeNet—as these layers capture rich, task-relevant representations without being biased by the final classification layer. For ResNet-18, we use the output of the final average pooling layer (i.e., the layer before the fully connected classifier). In DenseNet-121, we extract features from the global average pooling layer at the end of the final dense block. For LeNet-5, we use the output from the second fully connected layer just before the final softmax. These representations are aggregated per party and per window to form a latent representation per party per time window.
@@
Windowing Strategy. We adopt two different windowing strategies based on dataset characteristics. For large-scale datasets such as FMoW and Tiny-ImageNet-C, we use a tumbling window approach, where data in each window is disjoint and drawn from distinct covariate distributions—e.g., different timestamps in FMoW or corruption types in Tiny-ImageNet-C. This simulates strong distributional shifts between windows. For smaller datasets like Cifar-10-C, FEMNIST, and FashionMNIST, we use a sliding window approach, where overlapping samples are allowed across windows. This setup better supports gradual and more realistic distributional change in these datasets and allows us to simulate smoother transitions while controlling label and feature distributional shifts explicitly.
Metrics Captured: To evaluate model performance under distributional shifts, we focus on three complementary metrics: Accuracy Drop, Recovery Time, and Max Accuracy. Accuracy Drop measures the immediate decline in performance after a shift, indicating the model’s brittleness to sudden changes—larger drops imply greater sensitivity. Recovery Time captures the number of rounds required to regain 95% of pre-shift performance, reflecting the model’s adaptability and convergence speed. Max Accuracy denotes the highest accuracy achieved within a window, serving as a measure of the model’s generalization ability under the new distribution. Together, these metrics offer a comprehensive view of robustness, adaptability, and long-term effectiveness in non-stationary federated learning.
@@ 7. Results and Discussion
FedProx 20.39 44 63.48 39.57 ¿51 52.23 23.45 ¿51 45.07 18.24 ¿51 36.95
Fielding 21.32 ¿51 45.66 17.07 ¿51 48.29 15.76 ¿51 50.57 16.70 ¿51 48.34
OORT 8.60 ¿51 52.11 7.36 ¿51 50.80 4.68 ¿51 50.79 3.94 ¿51 50.06
ShiftEx 22.57 13 66.68 14.43 19 66.23 11.78 6 65.76 10.49 4 66.14
FedDrift 18.43 15 64.25 26.17 ¿51 62.33 20.84 ¿51 59.50 18.33 ¿51 54.13
@@
FedProx 17.80 ¿51 72.56 15.55 ¿51 77.55 17.01 ¿51 75.64 18.64 ¿51 76.35
Fielding 19.76 ¿51 72.25 16.81 ¿51 77.33 17.24 ¿51 75.48 15.39 ¿51 75.54
OORT 20.03 ¿51 72.24 16.66 ¿51 76.92 18.33 ¿51 75.37 16.39 ¿51 75.39
ShiftEx 37.51 ¿51 81.14 34.07 9 90.86 39.46 14 87.77 32.84 19 87.74
FedDrift 32.16 ¿51 78.47 27.01 18 86.95 33.15 18 84.67 29.73 23 83.68
@@
FedProx 27.19 ¿40 38.76 15.99 ¿40 37.37 11.49 ¿40 34.21 6.91 ¿40 40.01 11.43 ¿40 36.34
Fielding 29.93 ¿40 39.86 17.11 ¿40 38.42 12.40 ¿40 38.75 11.33 ¿40 43.61 15.05 ¿40 40.70
OORT 29.72 ¿40 54.76 18.87 ¿40 52.66 14.23 ¿40 52.30 10.23 ¿40 49.91 6.01 ¿40 50.83
ShiftEx 34.87 29 62.15 20.73 30 60.78 14.07 34 59.91 10.87 33 59.49 8.53 30 59.89
FedDrift 33.43 ¿40 52.94 20.79 ¿40 51.12 9.76 ¿40 53.27 10.99 ¿40 52.75 9.92 ¿40 51.36
@@
FedProx 9.36 ¿51 59.20 11.76 ¿51 55.26 2.24 ¿51 57.94 4.02 ¿51 56.38 1.38 ¿51 58.18
Fielding 9.14 48 60.18 11.12 ¿51 56.80 4.68 ¿51 57.86 4.72 ¿51 57.68 3.12 ¿51 58.24
OORT 6.76 ¿51 59.98 9.96 ¿51 54.44 4.40 ¿51 53.96 3.74 ¿51 54.78 0.26 ¿51 57.56
ShiftEx 8.90 42 62.72 10.62 35 61.58 4.32 14 63.56 3.94 1 64.98 3.12 0 66.36
FedDrift 8.14 40 62.68 15.04 50 60.40 10.56 44 60.68 9.82 ¿51 59.12 8.02 49 60.46
@@
FedProx 11.60 44 60.14 9.28 ¿51 59.68 5.96 ¿51 59.06 9.09 ¿51 59.42 9.33 ¿51 59.21
Fielding 8.68 30 62.16 11.08 47 60.64 5.14 ¿51 59.96 2.80 19 61.78 0.34 0 63.90
OORT 10.10 33 61.64 6.54 41 61.18 4.12 32 61.36 3.26 19 60.68 -0.86 0 63.04
ShiftEx 7.56 5 67.10 8.90 4 66.32 6.58 1 66.35 3.24 0 67.68 1.14 0 69.00
FedDrift 8.68 20 63.74 8.09 13 65.86 15.48 22 63.46 13.82 16 64.62 15.64 10 64.06

### v2

@@ Abstract.
Federated Learning (FL) enables collaborative model training across decentralized clients without sharing raw data, yet faces significant challenges in real-world settings where client data distributions evolve dynamically over time. This paper tackles the critical problem of covariate and label shifts in streaming FL environments, where non-stationary data distributions degrade model performance and necessitate a middleware layer that adapts FL to distributional shifts. We introduce ShiftEx  a shift-aware mixture of experts framework that dynamically creates and trains specialized global models in response to detected distribution shifts using Maximum Mean Discrepancy for covariate shifts. The framework employs a latent memory mechanism for expert reuse and implements facility location-based optimization to jointly minimize covariate mismatch, expert creation costs, and label imbalance. Through theoretical analysis and comprehensive experiments on benchmark datasets, we demonstrate 5.5-12.9 percentage point accuracy improvements and 22-95% faster adaptation compared to state-of-the-art FL baselines across diverse shift scenarios. The proposed approach offers a scalable, privacy-preserving middleware solution for FL systems operating in non-stationary, real-world conditions while minimizing communication and computational overhead.
@@
journalyear: 2025copyright: rightsretainedconference: 26th International Middleware Conference; December 15–19, 2025; Nashville, USAbooktitle: 26th International Middleware Conference (Middleware ’25), December 15–19, 2025, Nashville, USAdoi: 10.1145/3590140.3629123isbn: 979-8-4007-0177-1/23/12
@@ 1. Introduction
A critical challenge in streaming FL environments, is that the data from the parties will change and evolve over time. Consider the example shown in Figure 1 depicting satellite imagery to study land-use patterns. We observe differences in the images based on the changing weather conditions (e.g.) image clarity deteriorates in the winter when there is snow and frost. Hence, as new data arrives, both the input distribution $P(X)$ and label distribution $P(Y)$ can shift  (Zhao et al., 2018). These distributional shifts, known as covariate shift and label shift respectively, pose significant challenges for traditional FL approaches that maintain a single global model (Yoon et al., 2021b).
@@ 3.2. ShiftEx Architecture
Each client maintains a local dataset and runs a stream processing engine (e.g., Apache Kafka (Garg, 2013) or Flink (Carbone et al., 2015)) to collect, ingest, and preprocess incoming data streams, storing them in a local database. This setup resembles standard streaming data platforms. ShiftEx extends this setup by incorporating a shift detection module that computes latent feature representations and label distributions from local data across consecutive windows. Covariate shift is identified using Maximum Mean Discrepancy (MMD) on latent representations, while label shift is detected via Jensen–Shannon Divergence (JSD) on label histograms. After computing these statistics, clients transmit only the aggregated embeddings, label distributions, and corresponding MMD and JSD scores to the aggregator. MMD and JSD were selected because they are non-parametric and lightweight, introducing minimal overhead; however, the framework itself is detector-agnostic and can readily accommodate alternative choices if desired.
@@
A key distinction from centralized MoE systems is that expert assignment in ShiftEx occurs at the client level rather than at the granularity of individual inputs. Each client is associated with the most suitable expert (or, if necessary, a newly created one) based on latent statistics derived from its local data distribution. At inference time, the process is straightforward: the client simply uses the parameters of the expert model it was previously assigned and locally trained with. This avoids the need for a centralized gating network, ensures that inference is always well defined, preserves privacy by preventing raw inputs from being routed to a server, and reduces overhead by eliminating per-sample routing decisions.
@@ 4.2. Covariate Shift Detection via Maximum Mean Discrepancy
The key advantage of MMD is that it provides a principled way to detect distributional differences without requiring distributional assumptions or labeled data. It operates directly on feature representations, making it ideal for privacy-preserving covariate shift detection in federated settings. A high MMD value ($\delta_{\text{cov}}$) indicates a significant change in the feature distribution, signaling covariate shift.
@@ 5.2.2. Expert Assignment via Latent Memory Matching
Rather than training new experts for every shift, the aggregator maintains a latent memory, an exponential moving average of each expert’s embedding signatures—to enable expert reuse. For each covariate cluster $G_{j}$, the aggregator computes the cluster centroid $\bar{P}_{j}(X)$ and checks for matching existing experts:
$\text{If }\min_{k}\mathrm{MMD}(\bar{P}_{j}(X),\mathcal{M}^{(k)})\leq\epsilon\text{, then assign }G_{j}\text{ to expert }k$
where $\mathcal{M}^{(k)}$ represents the latent memory signature of expert $k$. This mechanism allows recurring covariate patterns to reuse existing experts without retraining, while preserving privacy since only aggregate embeddings are compared. It also provides a practical approximation to the MMD term in Equation 5.1, avoiding expensive pairwise comparisons across many clients. This process operationalizes the role of the $\lambda$ term in the optimization: $\lambda$ reflects the cost of spawning new experts, but in practice, no manual parameter tuning is required. Instead, we rely on clustering quality metrics, applying the Davies–Bouldin Index with the elbow method to determine when creating additional clusters (and thus new experts) is justified. In this way, the optimization captures the intuition of penalizing unnecessary expert proliferation, and the clustering criteria enforce this trade-off.
@@ 5.2.3. Expert Update using FLIPS/Small Groups
When a covariate cluster $G_{j}$ matches an existing expert $\theta^{(k)}$, the system assigns that expert to the cluster. Federated training is then performed using only the matched cohort, keeping expert specialization aligned with the current distribution. To avoid overfitting to skewed or imbalanced label distributions, we employ FLIPS (Federated Learning using Intelligent Participant Selection) (Bhope et al., 2023), which clusters clients by label proportions and selects participants equitably from each cluster. This ensures that the aggregated training data reflects a balanced label distribution and approximates the $\mu$ term in Equation 5.1. Here, $\mu$ encodes the penalty for label imbalance in the optimization objective; no manual tuning is required: FLIPS directly minimizes Jensen–Shannon Divergence across selected clients, preventing collapse of experts onto dominant classes. Thus, $\mu$ serves to formalize the intuition behind maintaining balanced label coverage, while the FLIPS mechanism realizes this balance automatically during training. For smaller covariate clusters ($|G_{j}|\leq\gamma$), no federated training is performed; instead, clients perform local finetuning or transfer learning on the assigned expert.
@@ 5.3. TEE for Privacy Preservation
FL ensures raw data never leaves client devices, it still exchanges intermediate artifacts such as model updates, embeddings, and drift statistics, which can leak private information via inference or reconstruction attacks (Hu et al., 2022; DLG git repository, 2020; Zhao et al., 2020; Geiping et al., 2020; Yin et al., 2021; Zhu et al., 2019). To strengthen end-to-end privacy, we augment ShiftEx with Trusted Execution Environments (TEEs) such as Intel SGX (int, [n. d.]) or AMD SEV (amd, [n. d.]), which provide hardware-level isolation and verifiable computation. In our design, clients compute $d$-dimensional embeddings locally and transmit them encrypted into a secure enclave, where drift detection (via MMD), clustering, and expert updates are performed without exposure to the aggregator. This prevents leakage of raw data, labels, or model parameters, while enabling secure aggregation of expert updates. TEEs incur modest overhead (e.g., 5% for AMD SEV) (Bhope et al., 2023) due to enclave initialization and memory constraints, but offer a practical trade-off between privacy and performance. Beyond TEEs, ShiftEx shares only aggregate statistics (MMD/JSD scores, label histograms), from which no known reconstruction attacks have been studied. Importantly, TEEs are optional: ShiftEx functions without them, but they provide a hardware-assisted layer of privacy protection.
@@ 6. Experimental Strategy
Datasets: We evaluate our framework on five datasets spanning diverse covariate and label shifts. FMoW (Functional Map of the World) (Christie et al., 2018) contains over a million satellite images labeled by functional use (e.g., hospital, airport, school) across regions and time; we select 10 labels (Jothimurugesan et al., 2023). Tiny-ImageNet-C (Hendrycks and Dietterich, 2019) augments Tiny-ImageNet with 15 corruption types at 5 severities, while CIFAR-10-C (Hendrycks and Dietterich, 2019) applies the same corruptions to CIFAR-10 for small-scale evaluation. FEMNIST (Caldas et al., 2019), a federated variant of EMNIST, includes handwritten characters from hundreds of users, and Fashion-MNIST (Xiao et al., 2017) consists of clothing images widely used as a benchmark. We simulate 200 parties for CIFAR-10-C, FEMNIST, and Fashion-MNIST to capture fine-grained heterogeneity in non-IID settings. For FMoW, we instead use 50 parties: its high-resolution imagery and naturally diverse covariates (geography, time, land use) already induce strong variability, so fewer parties suffice while ensuring each has enough samples for stable shift detection. In all cases, 20
@@
Models: We select architectures standard for each dataset: LeNet-5 (Lecun et al., 1998) for FEMNIST and FashionMNIST, DenseNet-121 (Huang et al., 2018) for FMoW, ResNet-18 (He et al., 2016) for CIFAR-10-C, and ResNet-50 (He et al., 2016) for Tiny-ImageNet-C. To detect covariate shifts, we extract encoder-based representations from the penultimate (pre-logit) layer of each model, which captures task-relevant features. Specifically, we use the final average pooling layer in ResNet-18, the global average pooling after the last dense block in DenseNet-121, and the second fully connected layer in LeNet-5. These embeddings are aggregated per party and per window to form latent representations.
@@
Windowing Strategy. We use two schemes matched to dataset scale and shift dynamics. For large datasets (FMoW, Tiny-ImageNet-C), we employ tumbling windows: each window is disjoint and drawn from distinct covariate distributions—e.g., different timestamps in FMoW and different corruption types in Tiny-ImageNet-C—thereby inducing strong between-window shifts. For smaller datasets (CIFAR-10-C, FEMNIST, Fashion-MNIST), we use sliding windows that allow overlap across windows to capture gradual change. FMoW windows align with natural temporal slices, whereas smaller datasets enforce a minimal stable party-size policy to ensure each party retains sufficient data for reliable training.
Metrics Captured: To evaluate model performance under distributional shifts, we focus on three complementary metrics: Accuracy Drop, Recovery Time, and Max Accuracy. Accuracy Drop measures the immediate decline in performance after a shift, indicating the model’s brittleness to sudden changes—larger drops imply greater sensitivity. Recovery Time captures the number of rounds required to regain 95% of pre-shift performance, reflecting the model’s adaptability and convergence speed. Max Accuracy denotes the highest accuracy achieved within a window, serving as a measure of the model’s generalization ability under the new distribution.
@@ 7. Results and Discussion
FedProx 20.39 $\pm$ 1.36 44 63.48 $\pm$ 1.38 39.57 $\pm$ 1.43 ¿51 52.23 $\pm$ 1.69 23.45 $\pm$ 2.47 ¿51 45.07 $\pm$ 1.79 18.24 $\pm$ 2.51 ¿51 36.95 $\pm$ 2.51
Fielding 21.32 $\pm$ 1.29 ¿51 45.66 $\pm$ 1.46 17.07 $\pm$ 1.83 ¿51 48.29 $\pm$ 1.57 15.76 $\pm$ 1.55 ¿51 50.57 $\pm$ 2.08 16.70 $\pm$ 2.18 ¿51 48.34 $\pm$ 2.18
OORT 8.60 $\pm$ 1.37 ¿51 52.11 $\pm$ 1.37 7.36 $\pm$ 1.39 ¿51 50.80 $\pm$ 1.85 4.68 $\pm$ 2.10 ¿51 50.79 $\pm$ 2.10 3.94 $\pm$ 2.70 ¿51 50.06 $\pm$ 2.70
ShiftEx 22.57 $\pm$ 0.64 13 66.68 $\pm$ 0.59 14.43 $\pm$ 0.76 19 66.23 $\pm$ 0.75 11.78 $\pm$ 0.91 6 65.76 $\pm$ 1.07 10.49 $\pm$ 1.09 4 66.14 $\pm$ 1.09
FedDrift 18.43 $\pm$ 1.37 15 64.25 $\pm$ 1.43 26.17 $\pm$ 1.63 ¿51 62.33 $\pm$ 1.70 20.84 $\pm$ 2.11 ¿51 59.50 $\pm$ 2.11 18.33 $\pm$ 1.89 ¿51 54.13 $\pm$ 1.89
@@
FedProx 17.80 $\pm$ 1.33 ¿51 72.56 $\pm$ 1.18 15.55 $\pm$ 1.48 ¿51 77.55 $\pm$ 1.81 17.01 $\pm$ 2.02 ¿51 75.64 $\pm$ 2.02 18.64 $\pm$ 2.14 ¿51 76.35 $\pm$ 1.96
Fielding 19.76 $\pm$ 1.48 ¿51 72.25 $\pm$ 1.72 16.81 $\pm$ 2.16 ¿51 77.33 $\pm$ 1.58 17.24 $\pm$ 2.03 ¿51 75.48 $\pm$ 2.03 15.39 $\pm$ 1.93 ¿51 75.54 $\pm$ 2.53
OORT 20.03 $\pm$ 1.36 ¿51 72.24 $\pm$ 1.28 16.66 $\pm$ 1.44 ¿51 76.92 $\pm$ 1.54 18.33 $\pm$ 2.06 ¿51 75.37 $\pm$ 2.06 16.39 $\pm$ 2.69 ¿51 75.39 $\pm$ 2.69
ShiftEx 37.51 $\pm$ 0.53 ¿51 81.14 $\pm$ 0.65 34.07 $\pm$ 0.83 9 90.86 $\pm$ 0.64 39.46 $\pm$ 1.02 14 87.77 $\pm$ 1.02 32.84 $\pm$ 1.09 19 87.74 $\pm$ 1.09
FedDrift 32.16 $\pm$ 1.60 ¿51 78.47 $\pm$ 1.20 27.01 $\pm$ 1.55 18 86.95 $\pm$ 1.63 33.15 $\pm$ 1.83 18 84.67 $\pm$ 1.83 29.73 $\pm$ 1.91 23 83.68 $\pm$ 1.91
@@
FedProx 27.19 $\pm$ 1.01 ¿40 38.76 $\pm$ 1.40 15.99 $\pm$ 1.53 ¿40 37.37 $\pm$ 1.44 11.49 $\pm$ 1.63 ¿40 34.21 $\pm$ 2.12 6.91 $\pm$ 2.76 ¿40 40.01 $\pm$ 2.16 11.43 $\pm$ 2.39 ¿40 36.34 $\pm$ 2.57
Fielding 29.93 $\pm$ 1.28 ¿40 39.86 $\pm$ 1.26 17.11 $\pm$ 1.40 ¿40 38.42 $\pm$ 1.52 12.40 $\pm$ 1.64 ¿40 38.75 $\pm$ 1.64 11.33 $\pm$ 1.91 ¿40 43.61 $\pm$ 2.42 15.05 $\pm$ 2.05 ¿40 40.70 $\pm$ 2.05
OORT 29.72 $\pm$ 1.13 ¿40 54.76 $\pm$ 1.18 18.87 $\pm$ 1.47 ¿40 52.66 $\pm$ 1.47 14.23 $\pm$ 2.12 ¿40 52.30 $\pm$ 2.12 10.23 $\pm$ 2.21 ¿40 49.91 $\pm$ 2.56 6.01 $\pm$ 2.56 ¿40 50.83 $\pm$ 2.56
ShiftEx 34.87 $\pm$ 0.52 29 62.15 $\pm$ 0.61 20.73 $\pm$ 0.69 30 60.78 $\pm$ 0.74 14.07 $\pm$ 0.91 34 59.91 $\pm$ 0.91 10.87 $\pm$ 0.99 33 59.49 $\pm$ 0.99 8.53 $\pm$ 1.09 30 59.89 $\pm$ 1.09
FedDrift 33.43 $\pm$ 1.46 ¿40 52.94 $\pm$ 1.36 20.79 $\pm$ 1.49 ¿40 51.12 $\pm$ 1.49 9.76 $\pm$ 1.61 ¿40 53.27 $\pm$ 1.61 10.99 $\pm$ 1.87 ¿40 52.75 $\pm$ 1.87 9.92 $\pm$ 2.21 ¿40 51.36 $\pm$ 2.21
@@
FedProx 9.36 $\pm$ 1.47 ¿51 59.20 $\pm$ 1.13 11.76 $\pm$ 1.19 ¿51 55.26 $\pm$ 1.53 2.24 $\pm$ 1.69 ¿51 57.94 $\pm$ 1.81 4.02 $\pm$ 1.96 ¿51 56.38 $\pm$ 1.96 1.38 $\pm$ 2.68 ¿51 58.18 $\pm$ 2.68
Fielding 9.14 $\pm$ 1.42 48 60.18 $\pm$ 1.06 11.12 $\pm$ 1.57 ¿51 56.80 $\pm$ 1.64 4.68 $\pm$ 1.92 ¿51 57.86 $\pm$ 1.92 4.72 $\pm$ 2.33 ¿51 57.68 $\pm$ 2.33 3.12 $\pm$ 2.59 ¿51 58.24 $\pm$ 2.59
OORT 6.76 $\pm$ 1.29 ¿51 59.98 $\pm$ 1.20 9.96 $\pm$ 1.45 ¿51 54.44 $\pm$ 1.40 4.40 $\pm$ 1.74 ¿51 53.96 $\pm$ 1.95 3.74 $\pm$ 2.19 ¿51 54.78 $\pm$ 2.60 0.26 $\pm$ 1.89 ¿51 57.56 $\pm$ 2.31
ShiftEx 8.90 $\pm$ 0.60 42 62.72 $\pm$ 0.71 10.62 $\pm$ 0.75 35 61.58 $\pm$ 0.73 4.32 $\pm$ 0.74 14 63.56 $\pm$ 0.75 3.94 $\pm$ 0.96 1 64.98 $\pm$ 0.84 3.12 $\pm$ 1.02 0 66.36 $\pm$ 1.02
FedDrift 8.14 $\pm$ 1.49 40 62.68 $\pm$ 1.58 15.04 $\pm$ 1.63 50 60.40 $\pm$ 1.41 10.56 $\pm$ 1.72 44 60.68 $\pm$ 1.72 9.82 $\pm$ 2.00 ¿51 59.12 $\pm$ 2.00 8.02 $\pm$ 2.30 49 60.46 $\pm$ 2.30
@@
FedProx 11.60 $\pm$ 1.15 44 60.14 $\pm$ 1.19 9.28 $\pm$ 1.80 ¿51 59.68 $\pm$ 1.55 5.96 $\pm$ 1.93 ¿51 59.06 $\pm$ 1.83 9.09 $\pm$ 2.32 ¿51 59.42 $\pm$ 2.32 9.33 $\pm$ 2.31 ¿51 59.21 $\pm$ 2.31
Fielding 8.68 $\pm$ 1.34 30 62.16 $\pm$ 1.29 11.08 $\pm$ 1.45 47 60.64 $\pm$ 1.45 5.14 $\pm$ 1.78 ¿51 59.96 $\pm$ 1.78 2.80 $\pm$ 2.41 19 61.78 $\pm$ 2.41 0.34 $\pm$ 2.70 0 63.90 $\pm$ 2.70
OORT 10.10 $\pm$ 1.12 33 61.64 $\pm$ 1.08 6.54 $\pm$ 1.75 41 61.18 $\pm$ 1.87 4.12 $\pm$ 1.78 32 61.36 $\pm$ 1.78 3.26 $\pm$ 1.83 19 60.68 $\pm$ 2.49 -0.86 $\pm$ 1.74 0 63.04 $\pm$ 2.70
ShiftEx 7.56 $\pm$ 0.56 5 67.10 $\pm$ 0.54 8.90 $\pm$ 0.75 4 66.32 $\pm$ 0.75 6.58 $\pm$ 0.74 1 66.35 $\pm$ 0.89 3.24 $\pm$ 0.76 0 67.68 $\pm$ 0.80 1.14 $\pm$ 1.03 0 69.00 $\pm$ 1.03
FedDrift 8.68 $\pm$ 1.29 20 63.74 $\pm$ 1.51 8.09 $\pm$ 1.73 13 65.86 $\pm$ 1.14 15.48 $\pm$ 1.50 22 63.46 $\pm$ 2.23 13.82 $\pm$ 2.50 16 64.62 $\pm$ 1.88 15.64 $\pm$ 2.51 10 64.06 $\pm$ 2.51